# Bayesian Concept Bottleneck Models with LLM Priors

### v2

@@ 1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [1, 2]. This has limited their adoption in high-stakes applications such as medicine [3] and settings where regulatory agencies require algorithms to be explainable [4]. Recent works have explored Concept Bottleneck Models (CBMs) [5, 6] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label. Thus, in principle, CBMs are much safer for use in high-stakes applications because humans can audit and, if necessary, modify the extracted concepts to fix predictions.
@@
Nevertheless, CBMs have yet to fully realize this promise. Early CBM methods were difficult to scale because they relied on human experts to specify and annotate concepts on training data. Recent works have proposed using LLMs to suggest and even annotate concepts [7, 8, 9, 10], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [9, 10]. This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [11]. Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [8], rather than “hard” binary values. However, “soft” CBMs can leak information about the label that would otherwise not have been available [12, 13, 14, 15, 16]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [15]. [^2] For instance, a human would fill in 0 when asked “does this bird have a gray breast” if there is no gray breast, but a “soft” CBMs may extract 0.2 if it sees a blue breast or gray head. This provides additional hints for downstream prediction. Because the definitions of soft concepts are ambiguous, humans can no longer easily interpret, audit, or intervene on such CBMs.
@@
The aim of this work is to minimize the accuracy gap without sacrificing the interpretability of the extracted concepts. To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [17, 10]. For instance, for the task of predicting hospital readmission, there are an infinite number of factors affecting a patient’s risk. Even a single concept like “smoking status” has an infinite number of refinements, such as the correlated concept of “whether a patient has quit smoking,” the broader concept of “substance use,” and more. As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [18].
@@
Rather than specifying concepts upfront, we investigate iterative refinement of concepts in a CBM with the assistance of an LLM. This updates a classical modeling paradigm for the LLM era: prediction models were traditionally designed through the collaboration between domain experts and data scientists, in which they iteratively refine/engineer features for simple tabular models. LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [19]. To set up guardrails, we “wrap” the LLM within a Bayesian posterior inference procedure to explore concepts with the help of an LLM in a statistically principled manner. This approach, which we refer to as Bayesian Concept bottleneck models with LLM priors (BC-LLM), offers the following benefits:
@@
  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$. In contrast, existing methods annotate hundreds or thousands of prespecified concepts [8].
@@
  • BC-LLM substantially improve the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models. Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
@@ 2 Related work
Prior CBM methods train on observations with hand-annotated concepts [5, 6, 20] and, more recently, concepts annotated by an LLM/VLM [21, 9, 10]. These methods are known to be highly sensitive to the selected set of concepts [18]. More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [7, 8, 22, 23], but this requires the LLM to have accurate prior knowledge about the supervised learning task. We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [12].
To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [17, 24]. These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM can typically only review a few observations due to limits on context length, so the resulting concept proposals are far from optimal.
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [25, 26, 27, 28, 29], which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [30].
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [31, 32, 33]. However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
Finally, recent works have considered combining LLMs with Bayesian techniques [34, 35, 36]. Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [37, 38], though recent work has proven that ICL is not fully consistent with Bayesian inference [39]. In contrast, this work studies how we can iteratively discover new concepts to include using LLMs using a Bayesian inference procedure, accounting for imperfections in LLM reasoning.
@@ 3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [40, 41, 42]. Here, we review the Bayesian solution for this more standard setting and then discuss how it may be extended to the CBM setting.
@@
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each outer loop, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [43]. Rather than immediately replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
@@
Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [37, 38], another idea is to include all the data when prompting the LLM to propose new concepts. That is, we ask the LLM itself to “conduct posterior inference” and interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from a conditional posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$ for some implicitly defined concept prior. This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [39]. Furthermore, we can no longer override poor concepts with data as the MH acceptance ratio simplifies to 1, i.e. the proposed concept is always accepted. So while Update 2 may generate concepts more efficiently, the sampling procedure may not converge to a good or even valid posterior distribution.
@@ 3.2 The Split-sample update: Propose with the LLM’s partial posterior
More specifically, the split-sample update first randomly splits the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). The LLM is prompted to generate candidate concepts by combining its prior knowledge with information from the data in $S$. We can then interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as the (conditional) partial posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [44]. This compares the likelihood of the held-out data $S^{c}$ with respect to the candidate and existing concepts—much like sample-splitting in frequentist settings—and thus provides an opportunity to correct inconsistencies between the LLM’s proposal and the actual posterior distribution.
@@
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [45]. Here, the LLM proposes a batch of candidates. We sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can prove that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition H.1 in the Appendix). Since LLMs are known to be imperfect Bayesian inference engines [39], this assumption is at best only approximately satisfied. Nevertheless, we can prove that the procedure still converges to the true concepts even when this assumption does not hold:
@@ 3.3 Putting it together: BC-LLM
Step 0: Initialization and keyphrase extraction We first use an LLM to extract “keyphrases” from each observation $x$, which will be used by the LLM to brainstorm candidate concepts. Here, “keyphrases” refer to short phrases: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be direct quotes or summarizations (e.g. diabetes) [46]. Keyphrases can be thought of as rough concepts, as they do not define a function but can help an LLM come up with a formal concept question (e.g. “diabetes” is a keyphrase whereas “Does the patient currently have diabetes?” is a concept). This brainstorming phase does not restrict the set of keyphrases the LLM can extract to allow for a wide exploration of concepts. The tradeoff is that the extracted keyphrases may be an incomplete summary of each observation. To initialize the CBM sampling procedure, these keyphrases are used to fit a (penalized) LR model for the target $Y$, similar to fitting a bag-of-words (BoW) model. An LLM then analyzes the most predictive keyphrases to generate an initial set of $K$ concepts and extracts these concepts from each observation.
@@
Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors
$p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ (1)
To estimate the integral (1), one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation for (1) [47], which is discussed in more detail in the Appendix. As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
Computational cost. BC-LLM performs $O(nTK)$ LLM queries, where $T$ is the number of outer loops. $K$ is typically chosen to be small, because a large $K$ yields less interpretable CBMs. Although standard Bayesian procedures typically draw thousands of posterior samples, we found that a small $T$ was quite effective in practice when combined with a greedy warm-start procedure; our experiments all use $T=5$. Generally speaking, a small $T$ suffices for achieving high prediction accuracy, a larger $T$ (e.g., 10) is helpful for quantifying the uncertainty of relevant concepts. Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [8]. In our experiments, BC-LLM was quick to run, as one iteration takes less than a minute for text data and a few minutes for image data on a normal laptop.

### v3

@@ 1 Introduction
Although machine learning (ML) algorithms have demonstrated remarkable predictive performance, many lack the interpretability and transparency necessary for human experts to audit their accuracy, fairness, and safety [dwork2012fairness, rudin2018please]. This has limited their adoption in high-stakes applications such as medicine [thirunavukarasu2023large] and settings where regulatory agencies require algorithms to be explainable [mesko2023imperative]. Recent works have explored Concept Bottleneck Models (CBMs) [koh2020concept, Kim2023-vu] as a potential solution: these methods leverage black-box algorithms to extract a small number of interpretable concepts, which are subsequently processed by a fully transparent tabular model to predict a target label. Thus, in principle, CBMs are much safer for use in high-stakes applications because humans can audit and, if necessary, modify the extracted concepts to fix predictions.
@@
Nevertheless, CBMs have yet to fully realize this promise. Early CBM methods were difficult to scale because they relied on human experts to specify and annotate concepts on training data. Recent works have proposed using LLMs to suggest and even annotate concepts [menon2022visual, Oikarinen2023-ee, McInerney2023-ab, Benara2024-lw], as LLMs are cheaper, often have sufficient world knowledge to hypothesize useful concepts, and can provide (relatively) high-quality concept annotations [McInerney2023-ab, Benara2024-lw]. This introduces new problems such as LLM hallucinations and inconsistencies but, more critically, does not resolve the fundamental problem that standard CBMs sacrifice a significant amount of accuracy to gain interpretability [yuksekgonul2022post]. Recent proposals try to address this by encoding concepts using “soft” continuous values, such as through embedding similarity [Oikarinen2023-ee], rather than “hard” binary values. However, “soft” CBMs can leak information about the label that would otherwise not have been available [mahinpei2021promises, margeloiu2021concept, havasi2022addressing, ragkousis2024tree, makonnen2025measuring]; this risk is particularly severe when concept extraction is jointly trained with the final label predictor [ragkousis2024tree]. [^2] For instance, a human would fill in 0 when asked “does this bird have a gray breast” if there is no gray breast, but a “soft” CBMs may extract 0.2 if it sees a blue breast or gray head. This provides additional hints for downstream prediction. Because the definitions of soft concepts are ambiguous, humans can no longer easily interpret, audit, or intervene on such CBMs.
@@
The aim of this work is to minimize the accuracy gap without sacrificing the interpretability of the extracted concepts. To this end, we revisit a known pain point of CBMs: Current training procedures require a pool of candidate concepts to be identified a priori, but ensuring this pool contains all relevant concepts is difficult, particularly in domains that are yet to be scientifically well-understood and where there are an infinite number of potentially relevant concepts [ludan2023interpretable, Benara2024-lw]. For instance, for the task of predicting hospital readmission, there are an infinite number of factors affecting a patient’s risk. Even a single concept like “smoking status” has an infinite number of refinements, such as the correlated concept of “whether a patient has quit smoking,” the broader concept of “substance use,” and more. As shown in prior work, when the candidate pool misses important concepts, the resulting CBMs may be inaccurate and even mislead users as to which concepts are truly relevant [Ramaswamy2023-qn].
@@
Rather than specifying concepts upfront, we investigate iterative refinement of concepts in a CBM with the assistance of an LLM. This updates a classical modeling paradigm for the LLM era: prediction models were traditionally designed through the collaboration between domain experts and data scientists, in which they iteratively refine/engineer features for simple tabular models. LLMs can help significantly accelerate this co-design process and scale up the number of iterations, but they are also imperfect query engines that can hallucinate, suggest bad candidate concepts, incorrectly annotate concepts, and may not even be self-consistent in their prior beliefs [Wang2023-ma]. To set up guardrails, we “wrap” the LLM within a Bayesian posterior inference procedure to explore concepts with the help of an LLM in a statistically principled manner. This approach, which we refer to as Bayesian Concept bottleneck models with LLM priors (BC-LLM), offers the following benefits:
@@
  • BC-LLM substantially improves the interpretability-accuracy tradeoff in real-world datasets: Experiments across multiple datasets and modalities (text, images, and tabular data) show that BC-LLM selects human-interpretable concepts while outperforming comparator methods, often even black-box models. Moreover, when assisting a hospital’s data science team to revise an existing tabular ML model, clinicians found BC-LLM to be substantially more interpretable and actionable.
@@
  • BC-LLM is cost-effective: BC-LLM selectively annotates $O(KT)$ concepts, where $K$ is the number of concepts in the CBM and $T$ is the number of outer loops; we often use only $K\approx 5$ and $T\approx 5$. In contrast, existing methods annotate hundreds or thousands of prespecified concepts [Oikarinen2023-ee].
@@ 2 Related work
Prior CBM methods train on observations with hand-annotated concepts [koh2020concept, Kim2023-vu, Yuksekgonul2023-um] and, more recently, concepts annotated by an LLM/VLM [Patel2023-dp, McInerney2023-ab, Benara2024-lw]. These methods are known to be highly sensitive to the selected set of concepts [Ramaswamy2023-qn]. More recent works have suggested using LLMs to prespecify the list of candidate concepts instead [menon2022visual, Oikarinen2023-ee, Yang2023-oh, panousis2023sparse, singh2025evaluating, xu2025graph, lam2024concept], but this requires the LLM to have accurate prior knowledge about the supervised learning task. We note that many of these works also make use of “soft concepts” and therefore suffer from information leakage [mahinpei2021promises, margeloiu2021concept, havasi2022addressing, ragkousis2024tree, makonnen2025measuring].
To find relevant concepts in a more data-driven manner, recent works have proposed taking an iterative approach [ludan2023interpretable, Liu2024-kq, sun2024general, kim2024constructing]. These methods employ a boosting-like approach, where the LLM is provided a set of misclassified examples and asked to add/revise concepts to improve performance. However, the LLM often struggles to identify relevant patterns from the presented examples, so the resulting concept proposals are far from optimal.
An alternative strategy is to conduct post-hoc distillation of a black-box model into interpretable concepts [Kim2018-ri, yeh2020completeness, Zhao2024-tj, Jiang2023-tx, singh2023augmenting], which can be used to learn an interpretable prediction model. Nevertheless, prior work has shown that the generated explanations are often not faithful to the black-box model, leading to poorer performance from the interpretable model [Doshi-Velez2017-uj].
Another related line of work uses LLMs to describe conceptual differences between pairs of datasets  [Zhong2023-rc, zhong2022describing, dunlap2024describing] or to directly describe a black-box function using natural-language concepts [kopf2025capturing, singh2023explaining, bills2023language, hoang2024llm]. However, existing methods are primarily suited for simple classification tasks, as they essentially learn a CBM with a single concept rather than a set of concepts.
Finally, recent works have considered combining LLMs with Bayesian techniques [yang2024bayesian, liu2024checkpoint, liu2024large]. Some have suggested that In-Context Learning (ICL) can be viewed as LLMs conducting Bayesian inference [Xie2021-gi, ye2024pre], though recent work has proven that ICL is not fully consistent with Bayesian inference [falck2024context]. In contrast, this work studies how we can iteratively discover new concepts to include using LLMs using a Bayesian inference procedure, accounting for imperfections in LLM reasoning.
@@ 3.1 Review: Bayesian variable selection
Learning CBMs shares many similarities with the more standard problem of learning sparse models in high-dimensional settings, where one must select among a finite number of features that have already been tabulated [ishwaran2008random, Castillo2015-pl, biswas2022scalable]. Here, we review the Bayesian solution for this more standard setting and then discuss how it may be extended to the CBM setting.
@@
Inference for $p(\vec{c}|\mathbf{y},\mathbf{X})$ is typically achieved through Gibbs sampling. For each outer loop, Gibbs rotates through indices $k=1,\ldots,K$, during which it replaces the $k$th concept in the current iterate $\vec{c}$ by drawing from the posterior distribution for $C_{k}$ conditional on the other concepts $\vec{c}_{-k}$. When one is unable to sample from the conditional posterior, Metropolis-within-Gibbs can be used instead [Gilks1995-fw]. Rather than immediately replacing the $k$th concept, it proposes a candidate concept $\check{c}_{k}$ given the other concepts per some distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ and accepts the candidate concept with probability $\alpha$. This acceptance probability $\alpha$, delineated in Algorithm 1, depends on the relative posterior probabilities of the candidate and existing concepts and is carefully designed so that the posterior distribution is stationary with respect to the sampling procedure.
@@
Update 2: Propose with the LLM’s “posterior.” Leveraging LLMs’ ability to perform ICL [Xie2021-gi, ye2024pre], another idea is to include all the data when prompting the LLM to propose new concepts. That is, we ask the LLM itself to “conduct posterior inference” and interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as sampling from a conditional posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y},\mathbf{X})$ for some implicitly defined concept prior. This approach is risky; while LLMs may be able to approximate posterior inference for small datasets, this approximation breaks down for larger sample sizes [falck2024context]. Furthermore, we can no longer override poor concepts with data as the MH acceptance ratio simplifies to 1, i.e. the proposed concept is always accepted. So while Update 2 may generate concepts more efficiently, the sampling procedure may not converge to a good or even valid posterior distribution.
@@ 3.2 The Split-sample update: Propose with the LLM’s partial posterior
More specifically, the split-sample update first randomly splits the data to construct subset $S$ of size $\lfloor\omega n\rfloor$ and its complement $S^{c}$, for some fraction $\omega$ (e.g. half). The LLM is prompted to generate candidate concepts by combining its prior knowledge with information from the data in $S$. We can then interpret the LLM’s proposal distribution $Q(C_{k};\vec{C}_{-k}=\vec{c}_{-k})$ as the (conditional) partial posterior distribution $p(C_{k}|\vec{c}_{-k},\mathbf{y}_{S},\mathbf{X})$. Then in the accept-reject step, the MH acceptance probability simplifies to the partial Bayes factor $\frac{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},(\vec{c}_{-k},\check{c}),\mathbf{X})}{p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})}$ [o1995fractional]. This compares the likelihood of the held-out data $S^{c}$ with respect to the candidate and existing concepts—much like sample-splitting in frequentist settings—and thus provides an opportunity to correct inconsistencies between the LLM’s proposal and the actual posterior distribution.
@@
In practice, rather than sampling a single candidate concept, we use a version of multiple-try Metropolis-Hastings (Algorithm 3) [liu2000multiple]. Here, the LLM proposes a batch of candidates. We sample a candidate from this batch based on their posterior probabilities and consider the sampled candidate for acceptance. This is more cost-effective as it lets us batch concept annotations, as described later.
Assuming that the LLM’s partial posterior across all iterations is consistent with some prior distribution $p(\vec{C})$, we can prove that the Markov chain defined by this procedure has the posterior $p(\vec{c}|\mathbf{y},\mathbf{X})$ as its stationary distribution (see Proposition H.1 in the Appendix). Since LLMs are known to be imperfect Bayesian inference engines [falck2024context], this assumption is at best only approximately satisfied. Nevertheless, we can prove that the procedure still converges to the true concepts even when this assumption does not hold:
@@ 3.3 Putting it together: BC-LLM
Step 0: Initialization and keyphrase extraction We first use an LLM to extract “keyphrases” from each observation $x$, which will be used by the LLM to brainstorm candidate concepts. Here, “keyphrases” refer to short phrases: keyphrases for images describe what’s in the image (e.g. blue eyes), while keyphrases for text data can be direct quotes or summarizations (e.g. diabetes) [Papagiannopoulou2020-vl]. Keyphrases can be thought of as rough concepts, as they do not define a function but can help an LLM come up with a formal concept question (e.g. “diabetes” is a keyphrase whereas “Does the patient currently have diabetes?” is a concept). This brainstorming phase does not restrict the set of keyphrases the LLM can extract to allow for a wide exploration of concepts. The tradeoff is that the extracted keyphrases may be an incomplete summary of each observation. To initialize the CBM sampling procedure, these keyphrases are used to fit a (penalized) LR model for the target $Y$, similar to fitting a bag-of-words (BoW) model. An LLM then analyzes the most predictive keyphrases to generate an initial set of $K$ concepts and extracts these concepts from each observation.
@@
Step 4: Accept/reject step The final step is to compute the sampling probabilities for the candidate concepts and the acceptance ratio, which involve comparing the proposal probabilities for the existing and candidate concept sets $\vec{c}$ as well as their split-sample posteriors $p(\mathbf{y}_{S^{c}}|\mathbf{y}_{S},\vec{c},\mathbf{X})=\int p(\mathbf{y}_{S^{c}}|\vec{\theta},\vec{c},\mathbf{X})p(\vec{\theta}|\mathbf{y}_{S},\vec{c},\mathbf{X})d\vec{\theta}.$ To estimate the integral, one approach is to sample from the posterior on $\vec{\theta}$, but this can be computationally expensive since one would need to implement posterior sampling in inner and outer loops. Alternatively, in the case of LR-based CBMs, one can use a Laplace-like approximation [Lewis1997-wl] (see Appendix). As the posterior distribution of LR parameters are asymptotically normal, this approximation becomes increasingly accurate as $n\rightarrow\infty$.
Computational cost. BC-LLM performs $O(nTK)$ LLM queries, where $T$ is the number of outer loops. $K$ is typically chosen to be small, because a large $K$ yields less interpretable CBMs. Although standard Bayesian procedures typically draw thousands of posterior samples, we found that a small $T$ was quite effective in practice when combined with a greedy warm-start procedure; our experiments all use $T=5$. Generally speaking, a small $T$ suffices for achieving high prediction accuracy, a larger $T$ (e.g., 10) is helpful for quantifying the uncertainty of relevant concepts. Critically, running BC-LLM is significantly more cost-effective than standard CBMs, which require $O(nW)$ queries where $W$ is the number of pre-specified concepts and must be large (typically hundreds or thousands) to avoid missing relevant concepts [Oikarinen2023-ee]. In our experiments, BC-LLM was quick to run, as one iteration takes less than a few minutes on a normal laptop. [check timings a bit more carefully]